please answer question 53 and show all the


please answer question 53 and show all the work
Answer #1

53)
Capaciance of each capacitor = A*eo/d
From figure we can make out that there are 3 Capacitors in series which make the tuning Capacitor,
So
Ceq = 1/(1/C + 1/C + 1/C)
Ceq = C/3
Ceq = A*eo/3d